OpenStudy (anonymous):

Partially derive with respect to y, the following: (Py/6EI)(6Lx-3x^2-vy^2)

OpenStudy (anonymous):

Can you retype the question

OpenStudy (anonymous):

it appears bit messy

OpenStudy (anonymous):

I can...one moment please. I'll try to do it with the equation editor.

OpenStudy (anonymous):

thanks , I am not able to answer wht Py/6ei is ?

OpenStudy (anonymous):

\[\left( \frac{ \delta }{ \delta y } \right) \frac{ Py }{ 6EI }\left( 6Lx-3x ^{2} -vy ^{2}\right)\]

OpenStudy (anonymous):

I'm pretty sure you're to use the product rule and the Py/6EI will end up being 1 because of the y in the numerator.

OpenStudy (anonymous):

O yeah I am used to product rule, dont worry, I was wondering whether it is part of question, or is a question number

OpenStudy (anonymous):

The solution looks like: \[\frac{ P }{ 6EI }\left( 6Lx-3x ^{2}-3vy ^{2} \right)\]

OpenStudy (anonymous):

I cannot seem to figure out ho to get the coefficient 3 for the last term.

OpenStudy (anonymous):

*how

OpenStudy (anonymous):

anyways here is your solution P/6EI (6Lx-3x^2-3vy^2)+(Py/6EI)*(0-0-6vy)

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

I'm not sure how you got that @psi9epsilon.

OpenStudy (anonymous):

Could you please clarify?

OpenStudy (anonymous):

product rule

OpenStudy (anonymous):

if the book answer is different then pls check the question, I dont see any error in the solution, unless rearragement of terms has been done